# Triple and binary evolution pathways play equally significant roles in LISA double white dwarf formation

### v1

@@ 4.2 Population properties
Figure 10: Population properties of LISA-detectable double white dwarfs from triple systems compared to isolated binaries. Overall, triple systems produce more massive white dwarfs than isolated binaries. For more details, see Section 4.2.
@@
In this section, we describe the similarities and differences in the population properties of double white dwarfs formed from triple systems and binary systems. Figure 10 shows the distribution of population properties such as chirp mass, primary mass, eccentricity, and gravitational-wave frequency. The green color corresponds to the properties of double white dwarfs from triple systems while the purple color shows the properties of double white dwarfs from binary star systems. The shaded green and purple regions represent the properties of resolvable double white dwarfs from triple and binary star systems, respectively. The resolved double white dwarf population traces the features of the total double white dwarf population. It is also interesting to note that all double white dwarfs in our mock Milky Way with $f_{\rm gw}>2\times 10^{-3}$ are fully resolvable by LISA. Individually resolving any double white dwarf with $f_{\rm gw}<2\times 10^{-3}$ is more difficult for LISA due to the higher degree of overlap in this frequency range.
@@
Triple systems produce about $\sim 1.5$ times more LISA double white dwarf systems with primary masses greater than $0.9\,\mathrm{M_{\odot}}$ and $\sim 3.9$ times more super-Chandrasekhar mass double white dwarfs (binary white dwarfs in which total mass of both the white dwarfs combined exceeds $1.44\,\mathrm{M_{\odot}}$) than the isolated binary channel. In addition, triple systems produce $\sim 1.6$ times more extremely low-mass white dwarfs ($m<0.25\,\mathrm{M_{\odot}}$) that enter the LISA frequency bandwidth than isolated binaries. Isolated binaries do not produce binaries with a chirp mass greater than $0.9\,\mathrm{M_{\odot}}$. In contrast, triple systems produce $\sim 10^{3}$ binaries whose chirp mass is greater than 0.9 $\mathrm{M_{\odot}}$. There are two possible ways to form massive white dwarfs: 1) The system originates from a massive main-sequence progenitor that evolves to a massive white dwarf; 2) Two less massive progenitors merge to form a massive main sequence progenitor that subsequently forms a massive white dwarf. In our simulations, initial massive main-sequence progenitors are rare. Hence, it is difficult to form massive white dwarfs in isolated binaries. However, in triple systems, the two less massive progenitors in the inner binary can merge to form a massive progenitor. The merged star co-evolves with the former tertiary component to form a double white dwarf with a massive component.
@@ 4.3 Different types of double white dwarfs
From the MSE single-star evolution model, Helium (He) white dwarfs typically originate from binary interactions where the progenitor loses its envelope before helium ignition, with typical masses of $\lesssim 0.45\,$M. Carbon-Oxygen (CO) white dwarfs form from intermediate-mass stars that exhaust helium in their cores and expel their outer layers, resulting in masses between $\sim 0.45-1.1\,$M. Oxygen-Neon (ONe) white dwarfs arise from more massive stars that undergo carbon burning before shedding their envelopes, with typical masses of $\gtrsim 1.1\,$M.
@@
Figure 11: Types of LISA-detectable double white dwarfs formed from triple systems compared to isolated binaries. In our models, ONe-ONe double white dwarfs are produced exclusively in triple systems. Error bars represent scaled up the fractional Poisson error from the intrinsic population evolved with MSE.
@@ 4.4 Detectability of the third star
We enfranchise again that approximately $57\,\%$ of all double white dwarfs detectable by LISA will have a bound tertiary companion. The bound tertiary star can be found in various evolutionary stages, including as a main sequence star, a giant star, or a white dwarf. In the remaining $43\,\%$ of systems, the third star either became unbound or there was a merger of two stars, reducing the system from a triple to a binary (cf. Section 3.3). If the tertiary is retained, it can impart accelerations to the center of mass of the binary, leading to observable Doppler shifts in the gravitational-wave signals (e.g., Seto, 2008; Robson et al., 2018; Tamanini & Danielski, 2019).
In the context of LISA Galactic binaries in hierarchical triple systems, Robson et al. (2018) identifies three regimes, essentially governed by the ratio of the outer orbital period to the observation time: 1) When the outer period is much larger than the observation time the hierarchical orbit imparts an overall unobservable Doppler shift. 2) When the outer period is up to a factor ten larger than the observation time the influence of the companion can be detected. 3) When the outer period is shorter than or comparable to the observation time, the eccentricity and period of the hierarchical orbit can be inferred. Specifically, a tertiary companion leaves a detectable imprint in the gravitational-wave signal if the outer binary period satisfies:
$T_{2}\lesssim T_{\mathrm{lim}}=43.2\,\text{yr} \left(\frac{\rho}{10}\cdot\frac{m_{3}}{1.0M_{\odot}}\cdot\frac{f}{5\,\text{mHz}}\right)^{3/4}\left(\frac{m_{1}+m_{2}}{2{\rm M}_{\odot}}\right)^{-1/2}$ (9)
×$\times\left(\frac{T_{\text{obs}}}{4\,\text{yr}}\right)^{3/8}\left(\frac{1+\frac{1}{2}e_{2}^{2}}{(1-e_{2}^{2})^{5/2}}\right)^{3/8},$
@@
[…]
Figure 12: Comparison between the outer orbital period of a triple system and the factor on the right-hand side of Eq. (9). Red points represents $T_{lim}$ vs outer orbital period $T_{2}$ for all the Galactic LISA double white dwarfs with a bound third star. The black line represents the points where $T_{2}=\text{factor}$. No tertiary star satisfies Eq. (9) for detectability within the LISA frequency bandwidth.
@@ 5.1 Comparison to previous works
We calculated the number and population properties of LISA double white dwarfs that originated from triple systems (cf. Section 2). We also simulated LISA double white dwarfs from an isolated binary population (cf. Appendix A) to compare with triple populations. Our isolated binary simulation predicts $\approx 10^{4}$ individually resolvable double white dwarfs in the LISA frequency bandwidth, which agrees with previous works, including Nelemans et al. (2004); Ruiter et al. (2010); Yu & Jeffery (2010); Korol et al. (2017); Lamberts et al. (2019); Li et al. (2023); Thiele et al. (2023); Tang et al. (2024). In total, our models predict $\sim 1.1\times 10^{7}$ double white dwarf sources that emit gravitational waves in the LISA frequency bandwidth but have very low $\rho$ to be individually detected by LISA.
@@
In particular, Korol et al. (2017) used binary population model of Toonen et al. (2012) based on the SeBa binary population synthesis code (Portegies Zwart & Verbunt, 1996) and an analytic Galactic potential and star formation history to estimate the number of Galactic double white dwarfs. They predict $\sim 2.6\times 10^{7}$ LISA double white dwarfs as foreground noise and $\sim 2.5\times 10^{4}$ as individually resolvable LISA double dwarf sources. Korol et al. (2022) performed a data-driven analysis using existing observational double white dwarf data and also estimated a LISA double white dwarf population of $\sim 2.6\times 10^{7}$ as foreground noise and $\sim 6.0\times 10^{4}$ as individually resolvable LISA double dwarf sources. Using the BSE code (i.e. same as out isolated binary evolution channel but with different underlying assumptions) combined with the Fire cosmological simulation, Lamberts et al. (2019) constructed the Galactic LISA double white dwarf population and estimated $\sim 6.2\times 10^{7}$ double white dwarf sources as foreground noise and $\sim 1.2\times 10^{4}$ as individually resolvable LISA double dwarf sources. Further, Li et al. (2023) also used the BSE code but with a mass transfer stability criterion by adopting critical mass ratios from the adiabatic mass loss model and estimated a foreground LISA double white dwarf population size of $\sim 5.0\times 10^{7}$ and about $4.0\times 10^{4}$ individually resolvable double white dwarfs. Our results are broadly consistent with previous studies in terms of the number of resolved binaries, but we estimate comparatively fewer double white dwarf sources emitting gravitational waves in the LISA frequency band (see Section 4.1.1 for details). This discrepancy may arise from multiple factors, including the use of different population synthesis codes, varying assumptions underlying binary evolution, and different approaches to modeling the Galaxy. Most importantly, previous studies model isolated binary evolution only, ignoring triple evolution channel.
@@
When comparing our isolated binary evolution results (with triple fraction set to zero, see Table 1) to previous studies, several factors can contribute to differences in LISA predictions. For instance, assuming a different total stellar mass for the Milky Way would linearly scale our results (cf. Eq. 2). As shown by Keim et al. (2023), the underlying star formation history also plays a role (see also Yu & Jeffery, 2010). However, the most likely primary source of differences is the modeling of stellar and binary evolution. This was recently demonstrated by van Zeist et al. (2024), who investigated the gravitational-wave population of the Large and Small Magellanic Clouds as a case study using the BPASS (Eldridge et al., 2017) and SeBa codes. They specifically attributed variations in the predicted number of LISA double white dwarfs to differences in the treatment of CE evolution and the stability of mass transfer. Indeed, even studies using the same population synthesis code report significant differences in LISA predictions when these processes are modeled differently (e.g., Korol et al., 2017; Li et al., 2023). As mentioned before, quantifying the impact of all these factors on LISA predictions is a large collaborative effort within the LISA Consortium’s Astrophysics Working Group. We refer the reader to those forthcoming results.
5.2 Uncertainties in our modelling
@@
Figure 13: Fractional difference of LISA-detectable double white dwarfs from models with varying common envelope efficiency, $\alpha_{\mathrm{CE}}$, and sub-solar metallicity, relative to the default values ($\alpha_{\mathrm{CE}}=1$, $Z=0.02$). The number of double white dwarfs varies by up to $8\,\%$ when the common envelope parameters are modified. Error bars represent scaled up the fractional Poisson error from the intrinsic population evolved with MSE.
@@
We use a critical mass ratio criterion to model the stability of mass transfer. Studies employing alternative stability criteria find that relatively fewer systems undergo a common envelope phase (Tauris et al., 2000; Podsiadlowski et al., 2002; Ge et al., 2010; Woods et al., 2011; Passy et al., 2012; Ge et al., 2015, 2020; Temmink et al., 2023; Li et al., 2023; van Zeist et al., 2024). Additionally, we assume that the orbit circularizes following the common envelope phase. However, there is ongoing debate about whether residual eccentricity may persist at the end of the common envelope evolution. The modeling of unstable mass transfer leading to a common envelope phase follows the approximate $\alpha$-$\lambda$ formalism. We investigated the effect of the common envelope efficiency parameter, $\alpha_{\text{CE}}$, on our results. For lower ($\alpha_{\text{CE}}=0.1$) and higher efficiency parameters ($\alpha_{\text{CE}}=10$), our simulations yield approximately $7.5\times 10^{6}$ and $7.6\times 10^{6}$ LISA double white dwarfs, respectively. For a lower value of the common envelope efficiency, the common envelope phase is relatively more effective at shrinking the orbit. Consequently, these systems enter the LISA frequency bandwidth within a Hubble time. In contrast, for the default value ($\alpha_{\text{CE}}=1$), the common envelope phase results in a relatively wider orbital period. However, for a high common envelope efficiency parameter, there is an increase in the number of LISA-detectable double white dwarf binaries. This is because the more efficient common envelope phase is less likely to merge binaries that would otherwise have merged for $\alpha_{\text{CE}}=1$.
@@
We assume that all progenitors of the Galactic LISA double white dwarfs are formed with solar metallicity. About $\sim 2\%$ of the star particles in the selected Milky Way model have subsolar metallicity. However, our Galaxy encompasses a larger range of metallicities. We also explore the impact of sub-solar metallicity by constructing the galaxy using the procedure described in Section 2.3 but with an initial metallicity of Z = 0.0002. It results in approximately $\sim 7.5\times 10^{6}$ LISA double white dwarfs, an increase of $\sim 0.3\times 10^{6}$ compared to those with solar metallicity. Stars with lower metallicity evolve relatively faster. For example, a $0.91\,\mathrm{M_{\odot}}$ star evolves into a white dwarf within a Hubble time at sub-solar metallicity, whereas it does not at solar metallicity. This increase in the number of low-mass stars that evolve into white dwarfs significantly contributes to the population of LISA-detectable double white dwarfs.
@@
The ZLK effect induces eccentric oscillations in the inner binary, increasing the possibility of mass transfer at periapsis or eccentric mass transfer. The MSE model treats eccentric mass transfer using an approximate prescription. Furthermore, MSE also uses approximate prescriptions for mass transfer from the third star onto the inner binary. In our simulation, about $5.5\,\%$ of Galactic LISA double white dwarfs undergo a TCE phase before reaching the LISA frequency bandwidth. Inputs from detailed hydrodynamical (Glanz & Perets, 2020) simulations of eccentric mass transfer and triple mass transfer are needed to improve the eccentric mass transfer and TCE prescriptions.
@@
We highlight that simulations of triple systems are computationally expensive. We simulated $10^{5}$ triple systems, producing $\sim 3\times 10^{3}$ LISA double white dwarfs. Deriving from the mass function in Figure 1, there is a very low probability of producing high-mass white dwarfs. Hence, uncertainties in the statistics on the number of white dwarfs increase with mass. There is also a possibility of producing ONe white dwarfs from 8-10 M stars. Our models are created from progenitors in the mass range 1-8 M. The sampling uncertainties in this mass range are too high to draw meaningful conclusions. In addition, we use random sampling to construct the initial triple population. However, more targeted sampling algorithms, such as STROOPWAFEL (Broekgaarden et al., 2019), could be employed to assess the impact of sampling uncertainties in the initial population on rare events.
@@
Finally, we note that MSE is a population synthesis code designed for statistical studies, and individual system modeling is not recommended. Due to variations in floating-point representation and rounding across different machines, the code may yield slightly different results when executed on different machines. However, these numerical difficulties average out on a large sample of the population. Given this inherent complexity, we limited the computational time to 10 hours per system. Within this time frame, $4\,\%$ systems did not complete their evolution. Of these, around $0.4\,\%$ were dynamically unstable, while the remaining systems were undergoing stable secular evolution. The majority of systems completed their evolution in under 2 hours. Additionally, only $0.1\,\%$ systems had white dwarfs or both components massive enough to evolve into white dwarfs within a Hubble time. These systems were included in error calculations, while the remaining incomplete systems were excluded without affecting the overall results.

### v2

@@ 4.2 Population properties
Figure 10: Population properties of LISA-detectable double white dwarfs from triple systems compared to isolated binaries. Triple-origin systems tend to produce more massive white dwarfs than isolated binaries. Outline (step) histograms represent the total LISA-detectable population, while filled histograms show the subset of systems individually resolvable with S/N ¿ 7. For more on the distributions shown, refer to Sect. 4.2. Green indicates triple-origin systems, and purple indicates isolated binaries. Thin and thick vertical lines show the 68.3% confidence intervals for the full and individually resolvable populations, respectively. The vertical gray lines indicate the full range of values spanned in each bin. Confidence intervals are estimated via bootstrap resampling and reflect the sampling uncertainty arising from the stochastic seeding of the Galaxy (see Appendix B for details).
@@
In this section we describe the similarities and differences in the population properties of double white dwarfs formed from triple systems and binary systems. Figure 10 shows the distribution of population properties such as chirp mass, primary mass, eccentricity, and gravitational wave frequency. The green color corresponds to the properties of double white dwarfs from triple systems while the purple color shows the properties of double white dwarfs from binary star systems. The shaded green and purple regions represent the properties of resolvable double white dwarfs from triple and binary star systems, respectively. The resolved double white dwarf population traces the features of the total double white dwarf population. It is also interesting to note that all double white dwarfs in our mock Milky Way with $f_{\rm gw}>2\times 10^{-3}$ are fully resolvable by LISA. Individually resolving any double white dwarf with $f_{\rm gw}<2\times 10^{-3}$ is more difficult for LISA due to the higher degree of overlap in this frequency range.
@@
Triple systems produce about $\sim 1.5$ times more LISA double white dwarf systems with primary masses greater than $0.9\,\mathrm{M_{\odot}}$ and $\sim 3.9$ times more super-Chandrasekhar mass double white dwarfs (binary white dwarfs in which total mass of both the white dwarfs combined exceeds $1.44\,\mathrm{M_{\odot}}$) than the isolated binary channel. In addition, triple systems produce $\sim 1.6$ times more extremely low-mass white dwarfs ($m<0.25\,\mathrm{M_{\odot}}$) that enter the LISA frequency bandwidth than isolated binaries. Isolated binaries do not produce binaries with a chirp mass greater than $0.9\,\mathrm{M_{\odot}}$. In contrast, triple systems produce $\sim 10^{3}$ binaries whose chirp mass is greater than 0.9 $\mathrm{M_{\odot}}$. There are two possible ways to form massive white dwarfs: 1) The system originates from a massive main-sequence progenitor that evolves to a massive white dwarf; 2) Two less massive progenitors merge to form a massive main-sequence progenitor that subsequently forms a massive white dwarf. In our simulations, initial massive main-sequence progenitors are rare. Hence, it is difficult to form massive white dwarfs in isolated binaries. However, in triple systems, the two less massive progenitors in the inner binary can merge to form a massive progenitor. The merged star co-evolves with the former tertiary component to form a double white dwarf with a massive component.
@@ 4.3 Different types of double white dwarfs
From the MSE single-star evolution model, helium-core white dwarfs originate from binary interactions where the progenitor loses its envelope before helium ignition, with typical masses of $\lesssim 0.45\,$M. Carbon-Oxygen core white dwarfs form from intermediate-mass stars that exhaust helium in their cores and expel their outer layers, resulting in masses between $\sim 0.45-1.1\,$M. Oxygen-Neon core white dwarfs arise from more massive stars that undergo carbon burning before shedding their envelopes, with typical masses of $\gtrsim 1.1\,$M.
@@
Figure 11: Types of LISA-detectable double white dwarfs formed from triple systems compared to isolated binaries. In our models, ONe-ONe double white dwarfs are produced exclusively in triple systems. Error bars represent a 68.3% confidence interval and are derived from bootstrap resampling and capture the statistical uncertainty introduced by the stochastic seeding of the Galaxy. (See Appendix B for details.)
@@ 4.4 Detectability of the third star
We emphasise again that approximately $57^{+8}_{-5}\,\%$ of all double white dwarfs detectable by LISA will have a bound tertiary companion. The bound tertiary star can be found in various evolutionary stages, including as a main-sequence star, a giant star, or a white dwarf. In the remaining $43\,\%$ of systems, the third star either became unbound or there was a merger of two stars, reducing the system from a triple to a binary (See Sect. 3.3). If the tertiary is retained, it can impart accelerations to the center of mass of the binary, leading to observable Doppler shifts in the gravitational wave signals (e.g., 2008ApJ...677L..55S; 2018PhRvD..98f4012R; 2019NatAs...3..858T).
In the context of LISA Galactic binaries in hierarchical triple systems, 2018PhRvD..98f4012R identifies three regimes, essentially governed by the ratio of the outer orbital period to the observation time: 1) When the outer period is much larger than the observation time the hierarchical orbit imparts an overall unobservable Doppler shift. 2) When the outer period is up to a factor ten larger than the observation time the influence of the companion can be detected. 3) When the outer period is shorter than or comparable to the observation time, the eccentricity and period of the hierarchical orbit can be inferred. Specifically, a tertiary companion leaves a detectable imprint in the gravitational wave signal if the outer binary period satisfies
[…]
$\times\left(\frac{T_{\text{obs}}}{4\,\text{yr}}\right)^{3/8}\left(\frac{1+\frac{1}{2}e_{2}^{2}}{(1-e_{2}^{2})^{5/2}}\right)^{3/8},$
@@
In our simulations, in all systems that have retained the tertiary companion, the outer orbits are too wide to have any detectable effect within the resulting LISA gravitational wave signal. Figure 12 compares the outer orbital period $\mathrm{T_{2}}$ with the factor on the right hand side of the Eq. (9). The black line shows the limit where $\mathrm{T_{2}}$ equals the factor on the right hand side. We find that none of our surviving triple systems lie within, or close to, the detectable limit. In addition, we find no systems where $T_{\text{orb}}\leq T_{\text{obs}}$ or $T_{\text{orb}}\approx 10\times T_{\text{obs}}$. The minimum outer semi-major axis across all systems is found to be approximately $21\,\text{au}$, which corresponds to an orbital period of around $63\,\rm yr$.
Figure 12: Comparison between the outer orbital period of a triple system and the factor on the right-hand side of Eq. (9). Red points represent $T_{\mathrm{lim}}$ vs. the outer orbital period $T_{2}$ for all the Galactic LISA double white dwarfs with a bound third star. The black line represents the points where $T_{2}=T_{\mathrm{lim}}$. No tertiary star satisfies Eq. (9) for detectability within the LISA frequency bandwidth.
@@ 5.1 Comparison to previous works
We calculated the number and population properties of LISA double white dwarfs that originated from triple systems (See Sect. 2). We also simulated LISA double white dwarfs from an isolated binary population (See Appendix A) to compare with triple populations. Our isolated binary simulation predicts $\approx 10^{4}$ individually resolvable double white dwarfs in the LISA frequency bandwidth, which agrees with previous works, including 2004MNRAS.349..181N; 2010ApJ...717.1006R; 2010A&A...521A..85Y; 2017MNRAS.470.1894K; 2019MNRAS.490.5888L; 2023A&A...669A..82L; 2023ApJ...945..162T; 2024arXiv240520484T. In total, our models predict $\sim 1.1\times 10^{7}$ double white dwarf sources that emit gravitational waves in the LISA frequency bandwidth but have too low $\rho$ to be individually detected by LISA.
@@
In particular, 2017MNRAS.470.1894K used the binary population model of too12 based on the SeBa binary population synthesis code (1996A&A...309..179P) and an analytic Galactic potential and star formation history to estimate the number of Galactic double white dwarfs. They predict $\sim 2.6\times 10^{7}$ LISA double white dwarfs as foreground noise and $\sim 2.5\times 10^{4}$ as individually resolvable LISA double dwarf sources. 2022MNRAS.511.5936K performed a data-driven analysis using existing observational double white dwarf data and also estimated a LISA double white dwarf population of $\sim 2.6\times 10^{7}$ as foreground noise and $\sim 6.0\times 10^{4}$ as individually resolvable LISA double dwarf sources. Using the BSE code (i.e., similar to our isolated binary evolution channel but with different underlying assumptions) combined with the Fire cosmological simulation, 2019MNRAS.490.5888L constructed the Galactic LISA double white dwarf population and estimated $\sim 6.2\times 10^{7}$ double white dwarf sources as foreground noise and $\sim 1.2\times 10^{4}$ as individually resolvable LISA double dwarf sources. Further, 2023A&A...669A..82L also used the BSE code but with a mass transfer stability criterion by adopting critical mass ratios from the adiabatic mass loss model by Ge_2010; Ge_2015; Ge_2020, and estimated a foreground LISA double white dwarf population size of $\sim 5.0\times 10^{7}$ and about $4.0\times 10^{4}$ individually resolvable double white dwarfs. Our results are broadly consistent with previous studies in terms of the number of resolved binaries, but we estimate comparatively fewer double white dwarf sources emitting gravitational waves in the LISA frequency band (see Sect. 4.1.1 for details). This discrepancy may arise from multiple factors, including the use of different population synthesis codes, varying assumptions underlying binary evolution, and different approaches to modeling the Galaxy. Most importantly, previous studies model isolated binary evolution only, ignoring triples.
@@
When comparing our isolated binary evolution results (with triple fraction set to zero; see Table 1) to previous studies, several factors can contribute to differences in LISA predictions. For instance, assuming a different total stellar mass for the Milky Way would linearly scale our results (See Eq. 2). As shown by 2023MNRAS.521.1088K, the underlying star formation history also plays a role (see also 2010A&A...521A..85Y). However, the most likely primary source of differences is the modeling of stellar and binary evolution. This was recently demonstrated by van_Zeist_2024, who investigated the gravitational wave population of the Large and Small Magellanic Clouds as a case study using the BPASS (2017PASA...34...58E) and SeBa codes. They specifically attributed variations in the predicted number of LISA double white dwarfs to differences in the treatment of CE evolution and the stability of mass transfer. Indeed, even studies using the same population synthesis code report significant differences in LISA predictions when these processes are modeled differently (e.g., 2017MNRAS.470.1894K; 2023A&A...669A..82L). As mentioned before, quantifying the impact of all these factors on LISA predictions is a large collaborative effort within the LISA Consortium’s Astrophysics Working Group. We refer the reader to those forthcoming results.
5.2 Uncertainties in our modeling
@@
Figure 13: Fractional difference of LISA-detectable double white dwarfs from models with varying common envelope efficiency, $\alpha_{\mathrm{CE}}$, and subsolar metallicity relative to the default values ($\alpha_{\mathrm{CE}}=1$, $Z=0.02$). The number of double white dwarfs varies by up to $8\,\%$ when the common envelope parameters are modified.
@@
We adopt the default critical mass-ratio criteria in MSE to model the stability of mass transfer. Studies investigating alternative stability criteria find that mass transfer tends to be more stable than we have assumed, suggesting that fewer systems undergo a common envelope phase (Tauris:2000toa; Podsiadlowski_2002; Ge_2010; Woods_2012; Passy_2012; Ge_2015; Ge_2020; Temmink_2023). This affects the formation of double white dwarfs (2012ApJ...744...12W), particularly in the LISA band (2023A&A...669A..82L; van_Zeist_2024). Additionally, we assume that the orbit circularizes following the common envelope phase. However, there is ongoing debate about whether residual eccentricity may persist at the end of the common envelope evolution. The modeling of unstable mass transfer leading to a common envelope phase follows the approximate $\alpha$-$\lambda$ formalism. We investigated the effect of the common-envelope efficiency parameter, $\alpha_{\text{CE}}$, on our results. As discussed in 2013A&ARv..21...59I, the commonly used parameter $\alpha_{\text{CE}}$ only describes an energetic “efficiency” under restrictive assumptions which are not known to be true. Contributions from sources of energy not included in the standard $\alpha$-$\lambda$ formalism may perhaps allow for an effective $\alpha_{\text{CE}}>1$. We would nonetheless be surprised if our model with effective $\alpha_{\text{CE}}=10$ is a good description of reality, but include that model to explore how sensitive our predictions are to extremely different outcomes of common-envelope phases.
For lower ($\alpha_{\text{CE}}=0.1$) and higher efficiency parameters ($\alpha_{\text{CE}}=10$), our simulations yield approximately $7.5\times 10^{6}$ and $7.6\times 10^{6}$ LISA double white dwarfs, respectively. For a lower value of the common envelope efficiency, the common envelope phase is more effective at shrinking the orbit. Consequently, these systems enter the LISA frequency bandwidth within a Hubble time. In contrast, for the default value ($\alpha_{\text{CE}}=1$), the common envelope phase results in a wider orbital period. However, for a high common envelope efficiency parameter, there is an increase in the number of LISA-detectable double white dwarf binaries. This is because the more efficient common envelope phase is less likely to merge binaries that would otherwise have merged for $\alpha_{\text{CE}}=1$.
@@
We assume that all progenitors of the Galactic LISA double white dwarfs are formed with solar metallicity. About $\sim 2\%$ of the star particles in the selected Milky Way model have subsolar metallicity. However, our Galaxy encompasses a larger range of metallicities. We also explore the impact of subsolar metallicity by constructing the galaxy using the procedure described in Sect. 2.3 but with an initial metallicity of Z = 0.002. It results in approximately $\sim 7.5\times 10^{6}$ LISA double white dwarfs, an increase of $\sim 0.3\times 10^{6}$ compared to those with solar metallicity. Stars with lower metallicity evolve faster. For example, a $0.91\,\mathrm{M_{\odot}}$ star evolves into a white dwarf within a Hubble time at subsolar metallicity, whereas it does not at solar metallicity. This increase in the number of low-mass stars that evolve into white dwarfs significantly contributes to the population of LISA-detectable double white dwarfs.
@@
The ZLK effect induces eccentric oscillations in the inner binary, increasing the possibility of mass transfer at periapsis or eccentric mass transfer. The MSE model treats eccentric mass transfer using an approximate prescription. Furthermore, MSE also uses approximate prescriptions for mass transfer from the third star onto the inner binary. In our simulation, about $5.5\,\%$ of Galactic LISA double white dwarfs undergo a TCE phase before reaching the LISA frequency bandwidth. Inputs from hydrodynamical (10.1093/mnras/staa3242) simulations of eccentric mass transfer and triple mass transfer are needed to improve the eccentric mass transfer and TCE prescriptions.
@@
We highlight that simulations of triple systems are computationally expensive. We simulated $10^{5}$ triple systems, producing $\sim 3\times 10^{3}$ LISA double white dwarfs. Deriving from the mass function in Fig. 1, there is a very low probability of producing high-mass white dwarfs. Hence, uncertainties in the statistics on the number of white dwarfs increase with mass. There is also a possibility of producing ONe white dwarfs from 8-10 M stars. Our models are created from progenitors in the mass range 1-8 M. The sampling uncertainties in this mass range are too high to draw meaningful conclusions. In addition, we use random sampling to construct the initial triple population. However, more targeted sampling algorithms, such as STROOPWAFEL (Broekgaarden_2019), could be employed to address the impact of sampling uncertainties in the initial population on rare events.
@@
Finally, we note that MSE is a population synthesis code designed for statistical studies, and individual system modeling is not recommended. Due to variations in floating-point representation and rounding across different machines, the code may yield slightly different results when executed on different machines. However, these numerical difficulties average out on a large sample of the population. Given this inherent complexity, we limited the computational time to 10 hours per system. The majority of systems completed their evolution in under 2 hours. Within the ten-hour limit, $4\,\%$ of systems did not complete their evolution. Of these, around $0.4\,\%$ were dynamically unstable, while the remaining systems were undergoing stable secular evolution. Only $0.1\,\%$ of the systems did not complete their evolution within ten hours and had white dwarfs or both components massive enough to evolve into white dwarfs within a Hubble time. This $0.1\,\%$ of the systems contribute to some uncertainty in our predictions for LISA double white dwarfs, while the remaining incomplete systems were excluded without affecting the overall results.